Help with a) and b)

a) Compute GlaxoSmithKlines weighted average cost of capital.

b) What is the NPV of the new machine and should GlaxoSmithKline replace the old machine with the new one?

GlaxoSmithKline plc is a pharmaceutical company. It is considering the replacement of one of its existing machines with a new model. The existing machine can be sold now for £8,000. The new machine costs £50,000 and will generate free cash flows of £11,416.55 pa. over the next 6 years. The corporate tax rate is 35%. The new machine has average risk. GlaxoSmithKline's debt-equity ratio is 0.5 and it plans to cost of equity is 13.10%
a) Compute GlaxoSmithKline's weighted average cost of capital.
b) What is the NPV of the new machine and should GlaxoSmithKline replace the old machine with the new one?
Answer #1

a) Weight of equity, we = 1 / (1 + D/E) = 1 / (1 + 0.5) = 67%, weight of debt, wd = 33%

=> WACC = wd x rd x (1 - tax) + we x re = 33% x 13.1% x (1 - 35%) + 67% x 13.10% = 10.00%

b) PV of cash flows of the new machine can be calculated using PV function

N = 6, PMT = 11,416.55, I/Y = 10%, FV = 0

=> Compute PV = $49,722.05

=> NPV = 49,722.05 - 50,000 + 8,000 = $7,722.05